9. How many 10-letter words are there in which each of the letters e, n, r , s occur

(a) At most once? (b) At least once?
Answer #1

This is a binomial distribution question with

n = 10

p = 4/26 = 0.1538

q = 1 - p = 0.8462


Please hit thumps up if the answer helped you